# ImageSet2Text: Describing Sets of Images through Text

### v1

@@ 3.2 Look and Keep
Next, all hypotheses in $\mathcal{H}$ are verified on the full image set $\mathcal{D}$ and the graph updated accordingly.
@@
ImageSet2Text evaluates each hypothesis $h_{i}$ in $\mathcal{H}$ against the entire image set $\mathcal{D}$ by leveraging the zero-shot classification capabilities of a CVL. This approach allows to set a one-vs-all classification problem, where positive and negative examples are generated for a given hypothesis $h_{i}$, drawing from $KG$. In particular, let $\mathcal{H}_{i}^{+}$ denote the set of alternative hypotheses that support $h_{i}$, which are constructed by substituting the object $o_{i}$ with its hyponyms in $KG$; and let $\mathcal{H}_{i}^{-}$ denote the set of alternatives that contradict $h_{i}$, which are constructed by substituting $o_{i}$ with its sibling nodes in $KG$. Note that the supporting set $\mathcal{H}_{i}^{+}$ is expanded to include $h_{i}$ itself. Next, the image set $\mathcal{D}$, along with the supporting set $\mathcal{H}_{i}^{+}$ and contradicting set $\mathcal{H}_{i}^{-}$, is projected into the CVL latent space, yielding the sets of embeddings $\mathcal{E}_{\mathcal{D}}$, $\mathcal{E}_{\mathcal{H}_{i}^{+}}$, and $\mathcal{E}_{\mathcal{H}_{i}^{-}}$, respectively. A $k$-Nearest Neighbors (kNN) classification algorithm is then applied using the embeddings from $\mathcal{E}_{\mathcal{H}_{i}^{+}}$ as positive examples and those from $\mathcal{E}_{\mathcal{H}_{i}^{-}}$ as negative examples, with cosine similarity serving as the weighting metric. In particular, the kNN classifies each image $x_{j}\in\mathcal{D}$ as supporting the hypothesis $h_{i}$ if its corresponding embedding $e_{j}\in\mathcal{E}_{\mathcal{D}}$ is labeled as positive, and as contradicting $h_{i}$ otherwise.
@@
As a result, the hypothesis $h_{i}$ is rejected if it is not verified on at least a predefined minimum portion $\alpha$ of the images in $\mathcal{D}$. Since hypotheses follow a hierarchical structure, if a hypothesis $h_{i}$ is rejected, then any more specific hypothesis $h_{i-1}\subset h_{i}$ is also invalid. This follows from the logical implication that $h_{i}\Rightarrow h_{i+1}$.
@@
Note that the hypotheses are verified in a general-to-specific manner to ensure semantic consistency in the CVL embedding space. If a highly specific hypothesis is prematurely tested without confirming its general category first, there is a risk of making comparisons in an embedding subspace that is not meaningful or reliable. Similarly, the sets $\mathcal{H}_{i}^{+}$ and $\mathcal{H}_{i}^{-}$ are used because computing cosine similarities in the embedding space without comparing them to any predefined reference does not directly provide a clear criterion to determine whether a hypothesis is valid or not [9].
Graph update.
At the end of the verification process, let $h_{*}=\langle s,p,o_{*}\rangle$ be the most specific hypothesis in $\mathcal{H}$ that has not been rejected. This hypothesis $h^{*}$ and its corresponding list of predicates $P$, which follow $o_{*}$, are then retained by appending them to the graph representation $\mathcal{G}_{\tau+1}$. Fig. 2 illustrates the updated graph at the end of the iteration.
@@ 3.3 Stopping Conditions
An iteration in ImageSet2Text interrupts if any of the following conditions occurs: 1) the VQA module flags a question as invalid (e.g., unsafe, inappropriate, unrelated to the content of the image) for at least a predefined number of images $\theta$ in $\mathcal{S}$; 2) no hypothesis in $\mathcal{H}$ is verified for $\mathcal{D}$; 3) the updated graph $\mathcal{G}_{\tau+1}$ adds no new information when compared to $\mathcal{G}_{\tau}$ as per an LLM evaluation.
The entire iterative process ends when: 1) no further graph expansion is possible, i.e., all existing nodes in $\mathcal{G}_{\tau}$ have been explored; or 2) a certain number $\epsilon$ of consecutive iterations are discarded according to the previously mentioned criteria.
Once the iterative process ends, any pending predicates is discarded, and the final description $d$ is generated directly from $\mathcal{G}=\mathcal{G}_{T}$ using the LLM, as illustrated in Fig. 2.
@@ 4 Evaluation
Describing large sets of images is a novel task, lacking suitable datasets and baselines for evaluation. We address this challenge by designing three evaluations that measure different aspects of the generated descriptions: (1) accuracy: i.e., alignment with visual content, (2) completeness: i.e., level of detail, and (3) readability and overall quality: i.e., human-evaluated coherence and ease of understanding.
Implementation details.
The implementation of ImageSet2Text reported in the experiments uses GPT-4o-mini [1] as the LLM; Open-CLIP ViT-bigG-14 [27] as the CVL; and WordNet [44] as the Knowledge Graph. The hyperparameters are set as: $M=10,\alpha=0.8,\theta=10,k=1,\delta=2,\epsilon=5$. Further implementation details are provided in App. A.
@@ 4.1 Accuracy
Accuracy is measured by how closely generated descriptions align with ground-truth descriptions of the image set. However, we are not aware of any publicly available benchmark dataset with ground-truth descriptions of large image sets. Therefore, as a proxy to evaluate the accuracy of the descriptions, we prompt an LLM to create a caption from the generated description of a set and leverage existing image captioning datasets to evaluate their accuracy.
Datasets.
We curate two datasets: (1) GroupConceptualCaptions, derived from Conceptual Captions [55] by grouping images with the same caption, using it as the group caption and (2) GroupWikiArt, derived from WikiArt [57] by grouping images with the same metadata (namely style, genre and artist) from which the group caption is constructed. This yields 221 image sets with $50$ to $4,112$ images each. Dataset details are in App. B.1.
The two proposed datasets pose different challenges. In GroupConceptualCaptions, the group caption can be generated from a single image since all images share the same caption. However, analyzing the entire group is essential for distinguishing meaningful features from irrelevant ones. In GroupWikiArt, individual images are less descriptive of the group, as the shared caption reflects abstract concepts, such as style and artist, that emerge only when examining a representative sample of images.
ImageSet2Text.
For each complete and detailed description generated by ImageSet2Text, we use an additional LLM call to generate a concise caption, extracting only the main elements of the description. Further details on this process are provided in App. B.3.
Baselines.
To the best of our knowledge, there are no publicly available baselines specifically designed for group image captioning at the scale explored in this work. Thus, we evaluate three established vision-language models for (single) image captioning: BLIP-2 [31], LLaVA-1.5 [36], and GPT-4V [1]. BLIP-2, optimized for image captioning and retrieval, serves as a strong baseline. LLaVA-1.5 enhances large language models with vision capabilities, enabling contextual and conversational image understanding. GPT-4V, as a state-of-the-art commercial model, [^2] Chatbot Arena LLM Leaderboard (Vision) https://lmarena.ai, Last Access: 3rd March 2025 offers advanced semantic comprehension across multiple images.
Since these baselines are designed to process only a single image at a time, we use three different approaches for creating group captions:
  • Grid: We create grids of varying sizes from a subset of images in $\mathcal{D}$ and prompt models to caption the grid. Results are reported for different grid sizes.
  • Average embedding: We input all images into the vision encoder of open-source models (BLIP-2 or LLaVA-1.5), average their embeddings to form a group embedding, and generate a caption of the average embedding using the language decoder.
  • Summary: We generate individual captions for all images and summarize them into a group caption with GPT-4.
Further details on baselines are available in App. B.2.
Metrics.
We rely on three types of metrics:
  • Model-free metrics: CIDEr-D [60], SPICE [3], METEOR [14] and ROUGE-L (F1) [34], that rely on lexical structure and token statistics to compare the generated captions with reference captions.
  • Model-based metrics: BERTScore (F1) [66] and LLM-as-a-judge [67], that leverage learned representations to capture semantic similarity [33, 37].
  • Reference-free metrics: namely CLIPScore [24], which measures the alignment between generated captions and images in the CLIP embedding space [51] without relying on reference captions.
More details on the metrics are provided in App. B.4.
Table 1: Results on GroupConceptualCaptions dataset. We consider CIDEr-D (D), SPICE (S), METEOR (M), ROUGE-L (R-L), BERTScore (BERT), LLM-as-a-judge (Judge) and CLIPScore (CLIP) as metrics. The best score is bold, second best underlined.
Model / Setting C S M R-L BERT Judge CLIP
LLaVA-1.5 1x1 grid 0.103 0.081 0.101 0.144 0.640 0.284 0.272
2x2 grid 0.046 0.102 0.079 0.106 0.619 0.181 0.266
3x3 grid 0.082 0.112 0.096 0.086 0.627 0.207 0.268
4x4 grid 0.092 0.121 0.110 0.089 0.632 0.198 0.273
Avg emb. 0.053 0.041 0.074 0.107 0.586 0.103 0.232
Summary 0.038 0.085 0.112 0.091 0.626 0.198 0.301
GPT-4V 1x1 grid 0.251 0.130 0.137 0.189 0.655 0.302 0.299
2x2 grid 0.120 0.084 0.110 0.098 0.623 0.155 0.297
3x3 grid 0.143 0.108 0.099 0.096 0.635 0.284 0.314
4x4 grid 0.146 0.105 0.104 0.099 0.649 0.276 0.315
Summary 0.132 0.104 0.096 0.104 0.631 0.129 0.314
ImageSet2Text 0.210 0.143 0.149 0.155 0.674 0.345 0.325
Results on GroupConceptualCaptions.
We compare ImageSet2Text to LLaVA-1.5 and GPT-4V on GroupConceptualCaptions. BLIP-2 is not eligible, as the original ConceptualCaptions [55] dataset was used to train this model. The results are provided in Tab. 1. Using only a single image randomly selected from the group (1x1 grid), works best for both LLaVA-1.5 and GPT-4V for model-free and model-based metrics overall. However, for the reference-free CLIPScore, larger grid sizes (4x4 grid) and summarizing individual captions as a group caption works best for the baseline methods.
ImageSet2Text outperforms the baselines on the model-based and reference-free metrics and also performs very competitively on the model-free metrics. The model-free metrics focus on explicit token overlap, measuring therefore if specific objects or actions have been identified. In contrast, the model-based metrics capture a more nuanced similarity between the generated caption and the reference caption [66, 67], aligning well with human judgement of caption quality. The strong performance of ImageSet2Text in this dataset highlights its ability to generate accurate descriptions and subsequently captions. Additional results for larger grid sizes are provided in App. B.5, which were omitted from Tab. 1 as they were always worse than the performance with small grid sizes.
Results on GroupWikiArt.
Results comparing BLIP-2, GPT-4V, and ImageSet2Text are provided in Tab. 2. As expected from how the group caption is created in this dataset, the baselines on a single image (1x1 grid), the average embedding, and summarizing individual captions, lead to poor performance. The baselines perform better with larger grid sizes, depending on the specific metrics. We observe that the captions created with ImageSet2Text are competitive on the model-free METEOR and ROUGE-L metrics. Furthermore, they perform very well under the model-based BERTScore and outperform all baselines on the LLM-as-a-judge metric.
Additional results for LLaVA-1.5 are provided in App. B.5, which we omitted in the main paper as it performed worse than BLIP-2. Note that CLIPScore is not a reliable metric to assess the accuracy of captions on this dataset because the groups are defined on a more contextual level (art history) that might go beyond visual features. Thus, we omit this metric from Tab. 2.
In sum, these results confirm that the captions derived from the descriptions generated with ImageSet2Text accurately describe the set of images, particularly when evaluated with model-based metrics that capture semantics, providing a good proxy for human judgment of accuracy.
Table 2: Results on GroupWikiArt dataset. We consider CIDEr-D (D), SPICE (S), METEOR (M), ROUGE-L (R-L), BERTScore (BERT) and LLM-as-a-judge (Judge) as evaluation metrics. The best score is bold, second best underlined.
Model / Setting C S M R-L BERT Judge
BLIP-2 1x1 grid 0.002 0.034 0.055 0.063 0.539 0.019
2x2 grid 0.046 0.060 0.060 0.091 0.542 0.038
3x3 grid 0.117 0.103 0.077 0.088 0.583 0.179
4x4 grid 0.092 0.102 0.070 0.076 0.577 0.132
5x5 grid 0.068 0.091 0.065 0.058 0.565 0.132
6x6 grid 0.052 0.079 0.047 0.047 0.551 0.085
7x7 grid 0.062 0.084 0.046 0.049 0.551 0.057
Avg emb. 0.004 0.054 0.076 0.086 0.576 0.028
Summary 0.082 0.032 0.069 0.075 0.574 0.085
GPT-4V 1x1 grid 0.002 0.012 0.064 0.064 0.558 0.028
2x2 grid 0.011 0.079 0.101 0.058 0.594 0.151
3x3 grid 0.023 0.117 0.116 0.050 0.613 0.208
4x4 grid 0.022 0.108 0.108 0.052 0.617 0.208
5x5 grid 0.021 0.140 0.108 0.053 0.621 0.160
6x6 grid 0.032 0.177 0.108 0.050 0.624 0.142
7x7 grid 0.025 0.159 0.106 0.045 0.623 0.113
Summary 0.003 0.028 0.049 0.036 0.560 0.075
ImageSet2Text 0.032 0.063 0.115 0.090 0.620 0.248
@@ 4.2 Completeness
The previous experiment assessed how accurately a caption obtained from the generated descriptions reflects the images in a given set. However, since this evaluation was based on captions obtained from the descriptions, it did not measure the completeness of the descriptions, i.e., how much detail they include. To evaluate completeness, we consider the downstream task of Set Difference Captioning (SDC) on the PairedImageSets (PIS) dataset [18].
PIS consists of $150$ image set pairs labeled with ground-truth differences, where the task is to identify the differences between the pairs of image sets. Each pair of sets is categorized by difficulty into easy, medium, and hard. To tackle this task, the PIS dataset authors introduced VisDiff [18], a proposer-ranker framework where an LLM-based proposer suggests potential differences between sets, and a ranker evaluates and ranks them through CLIP embeddings. As the starting point to generate potential differences, the proposer is given captions generated by BLIP-2 on subsets of the two original datasets.
We hypothesize that: if the descriptions created by ImageSet2Text were sufficiently complete and detailed, they would provide a stronger foundation for the proposer to identify dataset differences. Hence, we perform an experiment where ImageSet2Text generates descriptions of the pairs of image sets in the PIS dataset, namely sets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, independently. Then, the associated graph representations constructed for $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ serve as input to the VisDiff proposer-ranker framework. We compare this approach with VisDiff in Tab. 3. To generate this table, we relied on the same evaluation methodology and metric used by the authors of VisDiff [18].
Results.
Introducing the graph representations of the sets generated by ImageSet2Text improves the performance in all the cases except for Acc@5 on the easy sets, where the performance is the same as that of VisDiff, $0.99$. ImageSet2Text’s superior performance in the medium and hard cases suggest that the descriptions generated by ImageSet2Text are richer in meaningful details compared to the captions generated via BLIP-2. Furthermore, this approach accounts for some of the limitations reported by the authors of VisDiff. In App. C, we provide an analysis on specific failure cases of VisDiff [18].
Table 3: Results of the completeness evaluation on the PIS dataset [18]. The best performance is highlighted in bold.

### v2

@@ 3.2 Look and Keep
Next, all hypotheses $h_{i}\in\mathcal{H}$ are verified on the full image set $\mathcal{D}$ and the concept graph updated accordingly.
@@
ImageSet2Text evaluates each hypothesis $h_{i}$ $\in\mathcal{H}$ against the entire image set $\mathcal{D}$ by leveraging the zero-shot classification capabilities of a CVL, following a one-vs-all classification problem, where positive and negative examples are generated for a given hypothesis $h_{i}$, drawing from $\mathcal{G}_{l}$. Let $\mathcal{H}_{i}^{+}$ denote the set of other hypotheses that support $h_{i}$, which are constructed by substituting the object $o_{i}$ with its hyponyms in $\mathcal{G}_{l}$; and let $\mathcal{H}_{i}^{-}$ denote the set of hypotheses that contradict $h_{i}$, which are constructed by substituting $o_{i}$ with its sibling nodes in $\mathcal{G}_{l}$. Note that the supporting set $\mathcal{H}_{i}^{+}$ is expanded to include $h_{i}$ itself. In the given example, for $h_{1}=\langle\text{`{desert}'},\text{`{color}'},\text{`{yellow}'}\rangle$, supporting hypotheses replace ‘yellow’ with its hyponyms $o_{h}$, while contradicting ones use sibling nodes $o_{s}$:
[…]
@@
Next, all images in $\mathcal{D}$, along with each $o_{h}$ and $o_{s}$ are projected one-by-one into the CVL latent space and L2-normalized, yielding the sets of embeddings $\mathcal{E}_{\mathcal{D}}$, $\mathcal{E}_{\mathcal{H}_{i}^{+}}$, and $\mathcal{E}_{\mathcal{H}_{i}^{-}}$, respectively. A weighted $k$-Nearest Neighbors (kNN) classifier is then applied using the embeddings from $\mathcal{E}_{\mathcal{H}_{i}^{+}}$ as positive examples and those from $\mathcal{E}_{\mathcal{H}_{i}^{-}}$ as negative examples, with cosine similarity serving as the weighting metric. In particular, the kNN classifies each image $x_{j}\in\mathcal{D}$ as supporting the hypothesis $h_{i}$ if its corresponding embedding $e_{j}\in\mathcal{E}_{\mathcal{D}}$ is labeled as positive, and as contradicting $h_{i}$ otherwise. As a result, the hypothesis $h_{i}$ is rejected if it is not verified on at least a predefined minimum portion $\alpha$ of the images in $\mathcal{D}$. Since hypotheses follow a hierarchical structure, if a hypothesis $h_{i}$ is rejected, then any more specific hypothesis $h_{i-1}\subset h_{i}$ is also unverified. This follows from the logical implication that $h_{i}\Rightarrow h_{i+1}$.
@@
The hypotheses are verified in a general-to-specific manner to ensure semantic consistency in the CVL embedding space. If a highly specific hypothesis is prematurely tested without confirming its general category first, there is a risk of making comparisons in an embedding subspace that is not reliable (radford2021). Similarly, the set $\mathcal{H}_{i}^{-}$ is used because computing cosine similarities solely within $\mathcal{H}_{i}^{+}$ would not provide a reliable basis for evaluating the validity of a hypothesis. Comparisons against negative examples are necessary to establish a meaningful criterion for hypothesis acceptance (chattopadhyay2024).
Concept Graph update.
At the end of the verification process, let $h_{\checkmark}=\langle s,p,o_{\checkmark}\rangle$ be the most specific hypothesis $h_{i}\in\mathcal{H}$ that is verified; $h_{\checkmark}$ is appended to the graph $\mathcal{G}_{c}^{\tau+1}$ and the list of still‑pending predicates $P$ from $o_{0}$ is carried over to the next iteration (as illustrated in Fig. 3). The object $o_{\checkmark}$, now a leaf node in $\mathcal{G}_{c}^{\tau+1}$, is thus eligible to become the subject $s$ in a next iteration.
@@ 3.3 Stopping Conditions
An iteration in ImageSet2Text interrupts if any of the following conditions occurs: 1) the VQA module flags a question as invalid (e.g., unsafe, inappropriate, unrelated to the content of the image) for at least a predefined number of images $\theta$ in $\mathcal{S}$; 2) no hypothesis in $\mathcal{H}$ is verified for $\mathcal{D}$; 3) the updated graph $\mathcal{G}_{c}^{\tau+1}$ adds no new information when compared to $\mathcal{G}_{c}^{\tau}$ as per an LLM evaluation. The entire iterative process ends when: 1) no further graph expansion is possible, i.e., all existing nodes in $\mathcal{G}_{c}^{\tau}$ have been explored; or 2) a certain number $\epsilon$ of consecutive iterations are discarded according to the previously mentioned criteria. Once the iterative process ends, any pending predicate in $P$ is discarded, and the final textual description $d$ is generated directly from $\mathcal{G}_{c}=\mathcal{G}_{c}^{T}$ using the LLM, as illustrated in Fig. 3.
@@ 4 Evaluation
We report the results of evaluating the generated descriptions (qualitative examples available in App. B.2) according to three properties, as depicted in Fig. 2.
Implementation details
We use GPT-4o-mini (achiam2023) as the LLM (including for VQA), Open-CLIP ViT-bigG-14 (ilharco2021) as the CVL, and WordNet (miller1995) as the external lexical graph. Each random subset $\cal{S}$ contains $M=10$ images; hypotheses are rejected if their verification rate falls below $\alpha=0.8$ with $k=1$; iterations terminate upon encountering $\theta=10$ invalid images; traversing the knowledge hierarchy is limited to $\delta=2$ steps; and the maximum consecutive discarded iterations for stopping is $\epsilon=5$. Further details are provided in App. A.
@@ 4.1 Accuracy
Accuracy is evaluated through the task of group image captioning. This task tests whether models can identify and describe common visual elements in a group of images, making it an ideal proxy for this property of the descriptions.
Datasets
We did not find publicly available benchmarks for group image captioning of either small (up to 30 images) or large (up to thousands of images) image sets. Hence, we created two new benchmark datasets for this part of the evaluation (details in App. B.1):
  • GroupConceptualCaptions: Images with the same caption from the Conceptual Captions dataset (sharma2018) are grouped. Each group’s caption serves as ground truth, providing concise single-sentence descriptions of the main visual content. This dataset contains $116$ groups with a total of $23,412$ images.
  • GroupWikiArt: WikiArt artworks (artgan2018) with identical style, genre, and artist, are grouped, and the group captions are derived from these attributes. This dataset allows to evaluate nuanced, abstract concepts emerging from shared artistic interpretations within groups, and requires examining multiple images to capture group similarities. The dataset contains $105$ groups and $53,707$ images.
Models
Given the lack of public models for group image captioning, we compare ImageSet2Text with four state-of-the-art vision-language models: BLIP-2 (li2023), LLaVA-1.5 (liu2023), GPT-4o (achiam2023), and Qwen2.5-VL (bai:2025). As these are designed for single-image captioning, we adapt them to group captioning through three strategies: (1) prompting image grids of varying sizes in a single image, (2) averaging image embeddings before caption generation (for open-source models), and (3) summarizing individual captions into a group caption using GPT-4o. Further details are available in App. B.2 and App. B.7. Tests on multi-image settings are reported in App. B.6. For ImageSet2Text, we generate captions by summarizing the full description into a single short sentence using GPT-4o, as detailed in App. B.3.
Metrics
Accuracy is quantified by measuring the semantic alignment between generated descriptions and ground-truth captions on both GroupConceptualCaptions and GroupWikiArt. Results are reported as the average rank across seven standard captioning metrics: four model-free, [^2] CIDEr-D, SPICE, METEOR, ROUGE-L. two model-based, [^3] BERTScore F1, LLM-as-a-judge. and one reference-free. [^4] CLIPScore. Full evaluation details can be found in App. B.4 and B.5.
Figure 4: Accuracy as average rank across seven metrics on GroupConceptualCaptions (left) and GroupWikiArt (right). Only top ten methods shown; full results in App. B.5.
Results
As shown in Fig. 4, ImageSet2Text achieves the best performance on both GroupConceptualCaptions and GroupWikiArt, with average ranks of $2.50$ and $7.57$, respectively, outperforming all baselines. In contrast, the performance of the baselines is not consistent across datasets: while Qwen2.5-VL is the second-best model in GroupConceptualCaptions, BLIP-2, LLaVA-1.5, and GPT-4o exhibit stronger results in GroupWikiArt. Metric-wise analysis (detailed in App. B.5) reveals that ImageSet2Text performs particularly strong on model-based and reference-free metrics, which better capture semantic content and align more closely with human judgment (zhang2020).
@@ 4.2 Completeness
Completeness is evaluated by means of the Set Difference Captioning (SDC) task, which consists of identifying differences between two image sets, $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. To perform this comparison, we rely on the concept graphs generated by ImageSet2Text for each set. If the concepts are incomplete or vague, they would lead to failures in detecting differences between the sets. Hence, SDC serves as an effective proxy for evaluating the completeness of the underlying information used to generate the final descriptions—and, by extension, the completeness of the descriptions themselves.
Dataset
We adopt the methodology and dataset (PIS) described in (visdiff). The PIS dataset consists of $150$ contrastive set pairs (with $30,000$ images in total) spanning easy, medium, and hard distinctions. Consistent success in this task across varying difficult levels indicates that the description captures a range of relevant details necessary to differentiate $\mathcal{D}_{A}$ from $\mathcal{D}_{B}$.
Models
The PIS dataset was introduced together with VisDiff (visdiff), a proposer-ranker framework in which an LLM-based proposer suggests potential differences between image sets, and a ranker evaluates and ranks them using CLIP embeddings. The proposer relies on captions created with BLIP-2 from subsets of the two original datasets to identify differences between the two image sets. We compare VisDiff with ImageSet2Text by replacing the BLIP-2 captions with the concept graph representations produced by ImageSet2Text, keeping the rest of the framework unchanged (details in App. C.1 and C.3).
Metrics
We use standard metrics: acc@1 and acc@5.
Results
Results are shown in Table 1. Using ImageSet2Text improves the performance with respect to the original VisDiff in all metrics except for acc@5 on the easy sets, where both methods achieve near-perfect accuracy ($0.99$). Importantly, ImageSet2Text consistently improves accuracy on the medium and hard sets, showing that its richer semantic representations enable better identification of subtle differences between the image sets (e.g., distinguishing between cars with metallic paint and matte paint). Indeed, the lack of details in the BLIP-2 captions lead to a set of failure cases in VisDiff that are mostly addressed with ImageSet2Text (see App. C.2).
Table 1: Completeness evaluation on the PIS dataset. The best performance per difficulty category is shown in bold.